I need to complete the square:
2d^2=8+10d

Answer by Alan3354(69443)

I need to complete the square:
2d^2=8+10d
-----------
2d^2 - 10d - 8 = 0
d^2 - 5d - 4 = 0
The c term has to be (b/2)^2, or (-5/2)^2 = 25/4
So it's necessary to add 4+25/4 to both sides
d^2 - 5d + 25/4 = 41/4
Then take the sqrt
d - 5/2 = sqrt(41)/2 and -sqrt(41)/2
Add the 5/2
d = (5+sqrt(41))/2 and d = (5-sqrt(41))/2